Solve by completing the square m^2-49/2m=1/2?
Answer by stanbon(75887)

Solve by completing the square m^2-49/2m=1/2
-------------------------
m^2 - (49/2)m + (49/4)^2 = (32/16)+(49^2/16)
---------
(x-(49/4))^2 = 2433/16
-----------------------
x-(49/4) = sqrt(2433)/4 or x-(49/4) = -sqrt(2433)/4
========================================================
Cheers,
Stan H.
=============